Consider a simple economy with no government and no external sector. If the saving function S=-100+0.4Y, what is the value of the simple expenditure multiplier (write answer with only one decimal, e.g. 3.7)?
Answer #1

S = -100 + 0.4Y

Coefficient on Y is the marginal propensity to save

MPS = 0.4
Simple expenditure multiplier = 1/MPS = 1/0.4= 2.5